# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ B.1 Proofs of theoretical results for the Berkson error model
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]$ $=\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\|\mu_{\operatorname{\mathbb{P}}_{X}}-\mu_{\operatorname{\mathbb{F}}^{\prime}}\|_{\mathcal{H}_{k_{X}}}^{2}]$
$=\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}(\mu_{\operatorname{\mathbb{P}}^{i}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}})\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]$
$=\operatorname{\mathbb{E}}_{\xi^{1:n}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{1:n}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\left(\sum_{j=1}^{\infty}\xi_{j}^{i}k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right)\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]\right]$
$\leq\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left\|\sum_{j=1}^{\infty}\xi_{j}^{i}k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]\right]$
$=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left\|\sum_{j=1}^{\infty}\xi_{j}^{i}\left(k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right)\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]\right]$
$=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\sum_{j=1}^{\infty}(\xi_{j}^{i})^{2}\left\|k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right\|^{2}_{\mathcal{H}_{k_{X}}}\right.\right.$
$\quad\quad\left.\left.+2\sum_{j\neq t}\xi_{j}^{i}\xi_{t}^{i}\left<k_{X}(z_{j},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}},k_{X}(z_{t},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right>_{\operatorname{\mathcal{H}_{k_{X}}}}\right]\right].$ (32)
For any $j\neq t$ and each $i\in\{1,\dots,n\}$ we have:
@@
$=\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left<k_{X}(z_{j},\cdot)-\operatorname{\mathbb{E}}_{z_{j}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},\cdot)],k_{X}(z_{t},\cdot)-\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}^{\prime}}[(k_{X}(z_{t},\cdot)]\right>_{\operatorname{\mathcal{H}_{k_{X}}}}\right]$
@@
$\qquad-\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}_{w_{i}}}[k_{X}(z_{t},\cdot)]\operatorname{\mathbb{E}}_{z_{j}\sim\operatorname{\mathbb{F}}_{w_{i}}}[k_{X}(z_{j},\cdot)]+\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{t},\cdot)]\operatorname{\mathbb{E}}_{z_{j}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},\cdot)]$
$=\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]-\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]-\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]+\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]$
@@
$\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$ $=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)-\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}-2\left<k_{X}(z_{k},\cdot),\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}_{w_{i}}}[k_{X}(z_{k},\cdot)]\right>\right.$
$\quad\quad\left.+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]-2\left<\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)],\right.$
$\quad\quad\left.\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\right>+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]-2\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$\quad\quad+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}]-\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$\leq\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}]$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[|k_{X}(z_{k},z_{k})|].$ (34)
@@
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]$ $\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\operatorname{\mathbb{E}}_{\xi_{1:\infty}^{i}\sim\operatorname{\operatorname{GEM}}(c+1)}[(\xi_{j}^{i})^{2}]\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[|k_{X}(z_{k},z_{k})|]$
$\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\operatorname{\mathbb{E}}_{\xi_{1:\infty}^{i}\sim\operatorname{\operatorname{GEM}}(c+1)}[(\xi_{j}^{i})^{2}]$
$=\sum_{j=1}^{\infty}\frac{2(c+1)^{j-1}}{(c+3)^{j}(c+2)}$
$=\frac{2}{(c+3)(c+2)}\sum_{j=1}^{\infty}\left(\frac{c+1}{c+3}\right)^{j-1}$
$=\frac{2}{(c+3)(c+2)}\frac{1}{1-\frac{c+1}{c+3}}$
$=\frac{1}{c+2}$
@@
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]\leq\sqrt{\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]}\leq\frac{1}{\sqrt{c+2}}.$ (35)
@@
$\mu_{\operatorname{\mathbb{F}}^{\prime}}:=\mathbb{E}_{z\sim\operatorname{\mathbb{F}}^{\prime}}[k(z,\cdot)]=\frac{c}{c+1}\operatorname{\mathbb{E}}_{z\sim\operatorname{\mathbb{F}}}[k(z,\cdot)]+\frac{1}{c+1}\mathbb{E}_{z\sim\operatorname{\mathbb{P}}_{W}^{n}}[k(z,\cdot)]=\frac{c}{c+1}\mu_{\operatorname{\mathbb{F}}}+\frac{1}{c+1}\mu_{\operatorname{\mathbb{P}}_{W}^{n}}.$
@@
$\operatorname{\operatorname{MMD}}^{2}_{k_{X}^{i}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})$ $=\mathbb{E}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0},\nu^{\prime}\sim\operatorname{\mathbb{F}}_{N}^{0}}[k_{X}^{i}(\nu,\nu^{\prime})]-2\mathbb{E}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0},\nu^{\prime}\sim\operatorname{\mathbb{F}}_{N}}[k_{X}^{i}(\nu,\nu^{\prime})]+\mathbb{E}_{\nu\sim\operatorname{\mathbb{F}}_{N},\nu^{\prime}\sim\operatorname{\mathbb{F}}_{N}}[k_{X}^{i}(\nu,\nu^{\prime})]$
@@
and $\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})\leq\frac{1}{n}\sum_{i=1}^{n}\operatorname{\operatorname{MMD}}_{k_{X}^{i}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})=\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})$. Similarly,
@@
$\quad\quad\quad\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
@@
Lemma B.3 (Alquier and Gerber, (2024), Lemma 2).
For $\operatorname{\mathbb{P}}_{X}^{\prime},\operatorname{\mathbb{P}}_{X}^{\prime\prime}\in\mathcal{P}(\mathcal{X})$ let $P^{\prime}=P^{\prime}_{X}P_{Y\>|\>\cdot}$ and ′′$\operatorname{\mathbb{P}}^{\prime\prime}=\operatorname{\mathbb{P}}^{\prime\prime}_{X}\operatorname{\mathbb{P}}_{Y\>|\>\cdot}$. Assume that $(\operatorname{\mathbb{P}}_{Y\>|\>x})_{x\in\mathcal{X}}$ admits a bounded linear conditional mean embedding operator $\mathcal{C}_{Y\>|\>X}$. Then

### v3

@@ D.1 Proofs of theoretical results for the Berkson error model
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k_{X}}^{2}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]$
$\quad\quad=\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\|\mu_{\operatorname{\mathbb{P}}_{X}}-\mu_{\operatorname{\mathbb{F}}^{\prime}}\|_{\mathcal{H}_{k_{X}}}^{2}]$
$\quad\quad=\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}(\mu_{\operatorname{\mathbb{P}}^{i}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}})\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]$
$\quad\quad=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}\left[\|\mu_{\operatorname{\mathbb{P}}^{i}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\|^{2}_{\operatorname{\mathcal{H}_{k_{X}}}}\right]$
$\quad\quad\quad\quad+\frac{2}{n^{2}}\sum_{i\neq j}\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}\left[\left<\mu_{\operatorname{\mathbb{P}}^{i}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}},\mu_{\operatorname{\mathbb{P}}}^{j}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{j}}}\right>_{\mathcal{H}_{k_{X}}}\right]$
Notice that for each $i\neq j$, since $\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}^{i}\sim\mu_{i}}[\operatorname{\mathbb{P}}^{i}]=\operatorname{\mathbb{F}}^{\prime}_{w_{i}}$ by the definition of the mean of a DP, we have $\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\mu_{\operatorname{\mathbb{P}}^{i}}]=\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}$ and hence
$\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}\left[\left<\mu_{\operatorname{\mathbb{P}}^{i}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}},\mu_{\operatorname{\mathbb{P}}^{j}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{j}}}\right>_{\mathcal{H}_{k_{X}}}\right]$
$=\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\left<\mu_{\operatorname{\mathbb{P}}^{i}},\mu_{\operatorname{\mathbb{P}}^{j}}\right>]-2\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}\left[\left<\mu_{\operatorname{\mathbb{P}}^{i}},\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{j}}}\right>_{\mathcal{H}_{k_{X}}}\right]+\left<\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}},\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{j}}}\right>_{\mathcal{H}_{k_{X}}}$
$=0.$ (25)
Therefore, we have:
$\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k_{X}}^{2}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]$
$\quad\quad=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}\left[\|\mu_{\operatorname{\mathbb{P}}^{i}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\|^{2}_{\operatorname{\mathcal{H}_{k_{X}}}}\right]$
$\quad\quad=\frac{1}{n^{2}}\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left\|\sum_{j=1}^{\infty}\xi_{j}^{i}k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]\right]$
[…]
$\quad\quad\quad\quad\left.\left.+2\sum_{j\neq t}\xi_{j}^{i}\xi_{t}^{i}\left<k_{X}(z_{j},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}},k_{X}(z_{t},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right>_{\operatorname{\mathcal{H}_{k_{X}}}}\right]\right].$ (26)
For any $j\neq t$ and each $i\in\{1.\dots,n\}$ we have:
@@
[…]
$\left.\left.\quad\quad\quad k_{X}(z_{t},\cdot)-\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}^{\prime}}[(k_{X}(z_{t},\cdot)]\right>_{\operatorname{\mathcal{H}_{k_{X}}}}\right]$
@@
$\qquad-\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}_{w_{i}}}[k_{X}(z_{t},\cdot)]\operatorname{\mathbb{E}}_{z_{j}\sim\operatorname{\mathbb{F}}_{w_{i}}}[k_{X}(z_{j},\cdot)]$
$\quad\quad+\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{t},\cdot)]\operatorname{\mathbb{E}}_{z_{j}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},\cdot)]$
$=\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]-\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]$
$\quad\quad-\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]+\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]$
@@
$\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$
$\quad=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)-\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$
[…]
$\quad\quad\quad\left.+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$
$\quad=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]-2\left<\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)],\right$.
$\quad\quad\quad\left.\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\right>+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$\quad=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]-2\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$\quad\quad\quad+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$\quad=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}]-\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$\quad\leq\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}]$
$\quad=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[|k_{X}(z_{k},z_{k})|].$ (28)
@@
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k_{X}}^{2}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]$ $=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\operatorname{\mathbb{E}}_{\xi_{1:\infty}^{i}\sim\operatorname{\operatorname{GEM}}(c+1)}[(\xi_{j}^{i})^{2}]$
$\quad\quad\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[|k_{X}(z_{k},z_{k})|]$
$\leq\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\operatorname{\mathbb{E}}_{\xi_{1:\infty}^{i}\sim\operatorname{\operatorname{GEM}}(c+1)}[(\xi_{j}^{i})^{2}]$
$=\frac{1}{n}\sum_{j=1}^{\infty}\frac{2(c+1)^{j-1}}{(c+3)^{j}(c+2)}$
$=\frac{1}{n}\frac{2}{(c+3)(c+2)}\sum_{j=1}^{\infty}\left(\frac{c+1}{c+3}\right)^{j-1}$
$=\frac{1}{n}\frac{2}{(c+3)(c+2)}\frac{1}{1-\frac{c+1}{c+3}}$
$=\frac{1}{n(c+2)}$
@@
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]\leq\sqrt{\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k_{X}}^{2}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]}\leq\frac{1}{\sqrt{n(c+2)}}.$ (29)
@@
$\mu_{\operatorname{\mathbb{F}}^{\prime}}$ $:=\mathbb{E}_{z\sim\operatorname{\mathbb{F}}^{\prime}}[k(z,\cdot)]=\frac{c}{c+1}\operatorname{\mathbb{E}}_{z\sim\operatorname{\mathbb{F}}}[k(z,\cdot)]+\frac{1}{c+1}\mathbb{E}_{z\sim\operatorname{\mathbb{P}}_{W}^{n}}[k(z,\cdot)]$
$=\frac{c}{c+1}\mu_{\operatorname{\mathbb{F}}}+\frac{1}{c+1}\mu_{\operatorname{\mathbb{P}}_{W}^{n}}.$
@@
$\operatorname{\operatorname{MMD}}^{2}_{k_{X}^{i}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})$
$=\mathbb{E}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0},\nu^{\prime}\sim\operatorname{\mathbb{F}}_{N}^{0}}[k_{X}^{i}(\nu,\nu^{\prime})]-2\mathbb{E}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0},\nu^{\prime}\sim\operatorname{\mathbb{F}}_{N}}[k_{X}^{i}(\nu,\nu^{\prime})]+\mathbb{E}_{\nu\sim\operatorname{\mathbb{F}}_{N},\nu^{\prime}\sim\operatorname{\mathbb{F}}_{N}}[k_{X}^{i}(\nu,\nu^{\prime})]$
@@
and
$\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})\leq\frac{1}{n}\sum_{i=1}^{n}\operatorname{\operatorname{MMD}}_{k_{X}^{i}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})=\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N}).$
Similarly,
@@
$\quad\quad\quad\frac{1}{\sqrt{n(c+2)}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
@@
Lemma D.3 (Alquier and Gerber, (2024), Lemma 3).
For $\operatorname{\mathbb{P}}_{X}^{\prime},\operatorname{\mathbb{P}}_{X}^{\prime\prime}\in\mathcal{P}_{\mathcal{X}}$ let $\operatorname{\mathbb{P}}^{\prime}=\operatorname{\mathbb{P}}^{\prime}_{X}\times\operatorname{\mathbb{P}}_{Y\>|\>\cdot}$ and $\operatorname{\mathbb{P}}^{\prime\prime}=\operatorname{\mathbb{P}}^{\prime\prime}_{X}\times\operatorname{\mathbb{P}}_{Y\>|\>\cdot}$. Assume that $(\operatorname{\mathbb{P}}_{Y\>|\>x})_{x\in\mathcal{X}}$ admits a bounded linear conditional mean embedding operator $\mathcal{C}_{Y\>|\>X}$. Then